OpenStudy (anonymous):

on a right angle what is x when 25 and 25 are the other two numbers

OpenStudy (anonymous):

Is x the angle or the hypotenuse or what?

OpenStudy (anonymous):

If it is the hypotenuse then x^2 = (25^2 + 25^2) x = 1250^1/2 x = 25(2^1/2) x = 35.355

OpenStudy (anonymous):

You got a picture dude?